a plane flys 420 miles with the wind and 350 miles against the wind at the same length of time with a wind speed of 20 miles per hour what is the speed of the plane
Answer by mananth(16949)

plane speed x mph
wind speed 20 mph
..
against wind x-20 mph
with wind x+20 mph

Distance against wind 350
Distance with tail wind 420
Time = same
350/(x- 20)=420/(x+20)
350*(x+20)=420*(x-20)
350x+7000=420x -8400
420x-350x=8400+ 7000
70x=15400
/70
x= 220 mph